OpenStudy (anonymous):

Which of the following equations best represents the graph of f(p)? y = f(p) f(y) = p f(x) = p x = f(p)

OpenStudy (anonymous):

@Nnesha a?

OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value of x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

first is a yes :)

Nnesha (nnesha):

yep right 2nd one is right

OpenStudy (anonymous):

If f(x) = 2x2 - 30, find f(4). 35 -14 2 17

OpenStudy (anonymous):

c

OpenStudy (anonymous):

If f(x) is the total cost, in dollars, of x candies, which of the following statements best describes the meaning of f(2) = 6? The total cost of 2 candies is $6.00. The total cost of 6 candies is $2.00. The total cost of 2 candies is $3.00. The total cost of 3 candies is $2.00.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

x=f[p]

Nnesha (nnesha):

yep last two are also right

Nnesha (nnesha):

f(p) mean function of p which is same as y

Nnesha (nnesha):

so first one is wrong

OpenStudy (anonymous):

the very first one @Nnesha